OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 16? a) x squared over 16 b)±4square root of x c) ±square root of the quantity x plus 16 d) 1 over quantity x squared minus 16

OpenStudy (netlopes1):

Inverse function? Simply replace "x" with "y" and vice versa, like this:\[f(x)=x^2 - 16 \quad or \quad y=x^2 - 16 \rightarrow reversing \rightarrow x=y^2-16\]

OpenStudy (netlopes1):

Yet\[y^2=x+16 \rightarrow \quad Two \quad functions \rightarrow y=\sqrt{x+16} \quad or \quad y= - \sqrt{x+16}\]

OpenStudy (netlopes1):

Or simply\[y= \pm \sqrt{x+16}\], answer "c", ok?

OpenStudy (anonymous):

Thank you

OpenStudy (netlopes1):

you are welcome